# Towards universal property prediction in Cartesian space: TACE is all you need

### v1

@@ Abstract
Machine learning has revolutionized atomistic simulations and materials science, yet current approaches often depend on spherical-harmonic representations. Here we introduce the Tensor Atomic Cluster Expansion and Tensor Moment Potential, the first unified framework formulated entirely in Cartesian space for the systematic prediction of arbitrary structure-determined tensorial properties. TACE achieves this by decomposing atomic environments into a complete hierarchy of (irreducible) Cartesian tensors, ensuring symmetry-consistent representations that naturally encode invariance and equivariance constraints. Beyond geometry, TACE incorporates universal embeddings that flexibly integrate diverse attributes including basis sets, charges, magnetic moments and field perturbations. This allows explicit control over external invariants and equivariants in the prediction process. Long-range interactions are also accurately described through the Latent Ewald Summation module within the short-range approximation, providing a rigorous yet computationally efficient treatment of electrostatic interactions. We demonstrate that TACE attains accuracy, stability, and efficiency on par with or surpassing leading equivariant frameworks across finite molecules and extended materials, including in-domain and out-of-domain benchmarks, spectra, hessians, external-field response, charged systems, magnetic systems, multi-fidelity training, and heterogeneous catalytic systems. Crucially, TACE bridges scalar and tensorial modeling and establishes a Cartesian-space paradigm that unifies and extends beyond the design space of spherical-harmonic-based methods. This work lays the foundation for a new generation of universal atomistic machine learning models capable of systematically capturing the rich interplay of geometry, fields and material properties within a single coherent framework.
@@ 1 Introduction
In recent years, atomistic simulations and materials science have undergone a profound transformation. While Density Functional Theory (DFT) [1] remains the method of choice for balancing accuracy and generality, it is computationally demanding, with costs often scaling as $\mathcal{O}(N^{3})$ or higher. This significantly limits the range of systems and phenomena that can be realistically studied, restricting DFT’s applicability in areas such as high-throughput materials discovery, complex chemical reactions, and predictive modeling of large or dynamic systems. The growing availability of large, accurate first-principles datasets has consequently driven the development of machine learning interatomic potentials (MLIPs), which enable simulations at previously unattainable scales while delivering near ab initio accuracy across diverse chemistries.
MLIPs typically decompose the total energy into atomic site energy, which depend on the local environment of individual atoms. Early generations of MLIPs relied on fixed hand-crafted representations and simple Cartesian neural networks that mapped atomic coordinates and species to invariant features. These invariant approaches achieved considerable success by ensuring translational, rotational, and permutational symmetries [2, 3, 4, 5, 6]. However, they were fundamentally limited in accuracy and their ability to capture tensorial physical quantities.
The advent of equivariant message passing neural networks marked a major breakthrough. By incorporating irreducible representations of the O(3) group via spherical harmonics, models such as NequIP, Allegro and MACE [7, 8, 9] introduced physics-informed inductive biases that significantly improved data efficiency and generalization. Equivariant features within these networks transform consistently under rotations and inversions, allowing not only scalar properties such as energy to be predicted but also vectorial and tensorial properties such as dipole moments, and polarizabilities. This shift revolutionized the field and made spherical harmonic based methods the mainstream choice. Nevertheless, spherical tensor formulations present intrinsic challenges. The use of Clebsch-Gordan coefficients to couple angular momenta incurs substantial computational cost. Moreover, spherical tensors are defined with respect to a fixed axis, commonly the $z$ axis, which introduces directional bias [10, 11].
In parallel, Cartesian tensor approaches have experienced a resurgence as attractive alternatives. TensorNet [12], HotPP [13], CACE [14], and CAMP [15] explored feature construction directly in Cartesian space. These models rely on reducible Cartesian tensors and aimed to improve either accuracy or computational efficiency compared to spherical harmonic based methods. They indeed achieved some progress, particularly in low rank cases where Cartesian tensor contractions are generally faster. However, their accuracy has typically not surpassed that of models built on irreducible spherical tensor representations. TensorNet is restricted to rank-2 tensors and lacks systematic mechanisms for higher body interactions. HotPP [13] introduced deeper message passing and gating in Cartesian space but relied on reducible tensors with redundant representations. CACE encodes many body correlations explicitly but is limited in parameterization flexibility. CAMP [15] enforced strict permutation symmetry but did not guarantee tracelessness.
In fact, irreducible tensors also exist in Cartesian space. Unlike spherical tensors, they do not privilege any orientation, and their tensor products are often simpler and more computationally efficient up to rank four [11]. ICTP [11], although pioneering the use of irreducible Cartesian tensors to achieve accuracies comparable to spherical harmonics, requires explicit permutation of all indices in tensor products, and its implementation is limited to rank-3 tensors. Moreover, their representation cannot exploit lower-weight Cartesian tensors, thereby not fully realizing the potential of the Cartesian framework.
Beyond symmetry considerations, modern applications increasingly demand models that handle physical attributes beyond geometry, such as charge states, magnetic moments, and electronic temperature. Capturing these additional degrees of freedom requires a flexible and efficient tensorial framework capable of representing arbitrary combinations of physical quantities. Likewise, many properties of interest are responses to external perturbations, such as electric and magnetic fields, which require consistent handling of invariants and equivariants under external conditions. Long-range interactions pose an additional challenge. While models perform well in terms of scaling and accuracy under short-range approximations, they fail to effectively describe systems with significant long-range dependencies, where a rigorous treatment of electrostatics, such as Ewald summation, becomes essential. Most existing MLIPs do not integrate such long-range treatments seamlessly. Typically, model training is restricted to a single basis set or functional. We propose a multi-fidelity training strategy that simultaneously leverages datasets of varying accuracy, while also addressing several inherent limitations of the ACE framework.
Here we propose the Tensor Atomic Cluster Expansion (TACE), a new state-of-the-art framework that addresses these challenges by unifying scalar and tensorial modeling entirely within Cartesian space. We also introduce the Tensor Moment Potential (TMP) as a related architecture, however, the main focus of this work will be on TACE. TACE systematically decomposes atomic environments into a hierarchy of irreducible Cartesian tensors, ensuring exact symmetry consistency while avoiding the axis dependence and Clebsch-Gordan overhead of spherical harmonic based methods. By incorporating universal embeddings, TACE flexibly integrates diverse physical attributes including charges, magnetic moments, external fields and basis set into the representation, allowing explicit control over external invariants and equivariants in the prediction process. Crucially, TACE incorporates a latent Ewald summation (LES) module [16, 17, 18, 19], enabling a rigorous yet efficient treatment of long-range electrostatics within the short-range approximation. Beyond addressing the challenges of long-range interactions, we find that LES also resolves an inherent limitation of the ACE framework. Together, these innovations establish a Cartesian space paradigm that achieves accuracy, stability, and efficiency on par with or surpassing leading equivariant frameworks, while offering a systematic and extensible foundation for modeling the rich interplay of geometry, fields, and material properties.
@@ 2 Results
Models based on spherical harmonics equivariance have become widely adopted and extensively studied in recent years, providing a solid foundation for many applications in physics, chemistry, and machine learning. In contrast, Cartesian tensors, despite their natural and often more straightforward formulation in certain contexts, have not yet been presented in a systematic and comprehensive manner in the existing literature. To address this gap, we begin by reviewing the fundamental concepts of Cartesian tensors (2.1 and 2.2), which serve as the theoretical basis for our approach. We then proceed to describe our method in detail, highlighting its distinguishing features and advantages over existing frameworks.
@@ 2.1 Cartesian tensor
A generic Cartesian tensor of rank-$\nu$, denoted as $\prescript{\nu}{}{\mathbf{T}}$, is a multilinear object with $3^{\nu}$ components defined in a three-dimensional Euclidean space. The term generic Cartesian tensor refers to tensors without any imposed symmetry or traceless conditions among the indices. These components transform according to specific representation rules under the action of the orthogonal group $\mathrm{O(3)}$, which includes both rotations and spatial inversions in three-dimensional space. In general, such tensors are reducible under $\mathrm{O(3)}$, meaning that they can be decomposed into a direct sum of irreducible components, each associated with a definite angular momentum character $\ell$. A generic Cartesian tensor $\prescript{\nu}{}{\mathbf{T}}$ can be expressed as a sum of irreducible Cartesian tensors (ICTs) of the same rank but different weight $\ell$ (note that this “weight” is not the “learning weight” in machine learning):
@@
where each $\prescript{{(\nu;\ell;q)}}{}{\mathbf{T}}$ is an irreducible component of rank $\nu$, weight $l$ with $2\ell+1$ freedom, corresponding to a spherical tensor of degree $\ell$ and q denotes the label (since the multiplicity for given $\nu$ and $\ell$ may greater than 1). As a result, all $3^{\nu}$ components are independent, allowing the tensor to include all possible irreducible parts from weight $\ell=0$ up to $\ell=\nu$. This decomposition reveals that Cartesian tensors consist of a mixture of physical components: scalar parts ($\ell=0$), vector parts ($\ell=1$), and higher-rank fully symmetric-traceless tensors ($\ell\geq 2$), each transforming independently under rotation or inversion and corresponding to an invariant subspace of the tensor representation space.
@@ 2.2 Irreducible Cartesian tensor decomposition
As is well known, a generic rank-2 Cartesian tensor $\mathbf{T}$ can be decomposed into its irreducible components as follows:
${\mathbf{T}}_{ij}=\Bigg(\underbrace{\tfrac{1}{2}({\mathbf{T}}_{ij}+{\mathbf{T}}_{ji})-\tfrac{1}{3}\delta_{ij}{\mathbf{T}}_{kk}}_{\ell=2}\Bigg)+\Bigg(\underbrace{\tfrac{1}{2}{\mathbf{T}}_{ij}-{\mathbf{T}}_{ji})}_{\ell=1}\Bigg)+\Bigg(\underbrace{\tfrac{1}{3}\delta_{ij}{\mathbf{T}}_{kk}}_{\ell=0}\Bigg)$ (2)
This decomposition yields the rank-2 tensor’s irreducible components corresponding to angular momenta $\ell=2$, $\ell=1$, and $\ell=0$. Generalizations to higher-rank tensors with $\nu=3,4,5$ (numerical or analytical) were proposed in 1965, 1982, and 2024, respectively [20, 21, 22]. These ICTs could be obtained via irreducible Cartesian tensor decomposition (ICTD) matrices. We denote the ICTD operater and its associated decomposition matrix as $\mathscr{T}$ [23]. Although it is possible to compute ICTs with rank $\nu>5$ using numerical methods, such approaches are not suitable for use in the forward pass of deep learning models, as they significantly degrade computational efficiency and as far as we know, there are no models utilizing numerical methods for this operation now. In fact, decomposition matrices were previously only available up to rank $\nu=5$, and their generation has traditionally been computationally expensive. However, thanks to the recent work of ref. [23], analytical ICTD matrices for $\nu>5$ can now be efficiently constructed. The overall decomposition can be written as follows, where $\operatorname{vec}(\cdot)$ denotes flattening of the Cartesian tensor and the symbol $\triangleright$ denotes an operation applied to the operand on its right-hand side.
@@
2.3 The TACE architecture
Our feature learning begins with the construction of a graph in which each node is parameterized by learned ICTs ${\mathbf{h}}_{i}$ that encode its chemical environment. For each edge, a one-particle basis $\boldsymbol{\phi}_{ij}$ is computed and to obtain permutation-invariant representations of semi-local environments, a pooling operation is applied to aggregate these local features, resulting in the so-called atomic basis ${\mathbf{A}}_{i}$. Building on this, higher-body-order features are systematically constructed through tensor contraction over the ${\mathbf{A}}_{i}$, yielding a product basis ${\mathbf{B}}_{i}$ that captures ($n+1$)-body interactions. TACE follows the framework of Atomic Cluster Expansion (ACE) [24, 25, 9, 26, 27], while acknowledging that various alternative design choices remain available.
@@ 2.3.1 Equivariant operation
The first type, tensor contracion maps two tensors $^{\nu_{1}}\mathbf{T}$ and $^{\nu_{2}}\mathbf{T}$ to a new tensor $^{\nu_{3}}\mathbf{T}$ whose rank $\nu_{3}=\nu_{1}+\nu_{2}-2k$ is determined by the the number of contracted indices $k$, where $0\leq k\leq\min(\nu_{1},\nu_{2})$. When no indices are contracted, the operation is equivalent to tensor product. Eqn. 4 gives the element-wise expression of the tensor contraction, which may be abbreviated as $\prescript{\nu_{3}}{}{\mathbf{T}}=\prescript{\nu_{1}}{}{\mathbf{T}}\boldsymbol{\otimes}_{(\nu_{1},\nu_{2},\nu_{3})}\prescript{\nu_{2}}{}{\mathbf{T}}$. In practice, certain constants may be used to modulate the variance.
@@
We observe that for a given combination of tensor ranks $(\nu_{1},\nu_{2},\nu_{3})$, there can be multiple distinct ways to perform contractions, each corresponding to a different contraction path. Drawing inspiration from e3nn [28], we define each such distinct contraction configuration as a path. Specifically, if $k$ indices are contracted between the first and second tensors, the number of possible paths can be computed without considering any symmetries. It equals the number of ways to select $k$ contracted indices from the first tensor and $k$ contracted indices from the second tensor, multiplied by the number of permutations of these $k$ indices. Since Cartesian tensors correspond to reducible representations and therefore contain redundant information, Cartesian-based-tensor-contraction models [13, 15] typically include only a single path for each combination and may directly embedding radial information to edge. In contrast, by employing ICTD, we have the flexibility to incorporate either all possible paths or only a single path to achieve a desirable effect, and assign a radial-dependent weight to each.
The second type of operation consists of linear combinations of tensors with identical rank $\nu$, resulting in new tensors of the same rank. Specifically, given $\tilde{c}$ input tensors $\prescript{\nu}{\tilde{c}}{\mathbf{T}}$, we can construct $c$ output tensors $\prescript{\nu}{c}{\mathbf{T}}$ of the same rank through a learnable linear transformation. This operation can be interpreted as treating each tensor as a neuron, thereby enabling linear operations to be applied. In fact, similar to approaches such as NequIP and HotPP [7, 13], one could introduce nonlinearity via gating to construct multi-layer perceptrons (MLPs) for tensor with $\nu>0$. However, in TACE, we apply only linear operations to tensors with $\nu>0$, without introducing additional nonlinearities. Nonlinearity is introduced only in the radial network and in the final readout layer.

### v2

# TACE: A unified Irreducible Cartesian Tensor Framework for Atomistic Machine Learning

@@ Abstract
Machine learning has transformed atomistic simulations and materials science, yet most existing approaches rely on spherical-tensor representations. Here, we introduce the Tensor Atomic Cluster Expansion (TACE), a unified framework formulated entirely in Cartesian space, enabling systematic and consistent prediction of arbitrary structure-dependent tensorial properties. TACE achieves this by decomposing atomic environments into a complete hierarchy of irreducible Cartesian tensors, ensuring symmetry-consistent representations that naturally encode invariance and equivariance constraints. Beyond geometry, TACE incorporates universal embeddings that flexibly integrate diverse attributes including computational levels, charges, magnetic moments and field perturbations. This allows explicit control over external invariants and equivariants in the prediction process. Long-range interactions are also accurately described through the Latent Ewald Summation module within the short-range approximation, providing a rigorous yet computationally efficient treatment of electrostatic and dispersion effects. We demonstrate that TACE attains accuracy, stability, and efficiency on par with or surpassing leading equivariant frameworks across finite molecules and extended materials. This includes in-domain and out-of-domain benchmarks, spectra, Hessian, external-field responses, charged and magnetic systems, multi-fidelity training, heterogeneous catalysis, and even superior performance within the uMLIP benchmark. Crucially, TACE bridges scalar and tensorial modeling and establishes a Cartesian-space paradigm that unifies and extends beyond the design space of spherical-tensor-based methods. This work lays the foundation for a new generation of universal atomistic machine learning models capable of systematically capturing the rich interplay of geometry, fields and material properties within a single coherent framework.
@@ 1 Introduction
Atomistic simulations rely on accurate potential energy surfaces (PES) to describe structures, dynamics, and reactions across molecules and materials. While density functional theory (DFT) is the reliability reference, $\mathcal{O}(N^{3})$ or higher scaling limits the accessibility of atomistic simulations to large spatial and temporal scales and to complex chemical reactions, motivating machine learning interatomic potentials (MLIPs) as practical surrogates. Recent equivariant models have markedly advanced data efficiency and accuracy by learning from ab initio energies, forces, and stresses (E/F/S); yet can still yield limited fidelity in downstream physical property predictions despite low test errors when essential physical constraints are not included [1] or when the models show poor extrapolation and robustness [2]. Many physically meaningful quantities that are readily available from ab initio calculations, such as charges, magnetic moments, and external fields, are often underutilized or not effectively integrated into state-of-the-art (SOTA) models, which prevents one from realistic, cross-validated modeling and gleaning more insights.
Various physical constrains and conservation laws can provide critical information to assist the prediction of complex reaction systems. For quantities such as forces, which are explicitly derived from conservative field, it is essential to respect the underlying physical constraints to ensure a smoothly varying and physically meaningful landscape. In practice, direct-force learning (e.g., direct-force pretraining [1]) can be helpful, but robust and accurate prediction ultimately hinges on conservative training and physics-aware supervision. In electrocatalysis, applied potentials and external fields shift adsorption energies, stabilizing or destabilizing transition states, and altering reaction kinetics and spectroscopic responses; models that do not explicitly encode external fields and long-range interaction can fit forces yet yield physically incorrect conclusions. Likewise, magnetic materials are intrinsically 6N-dimensional rather than 3N (see SI Fig. S1, so restricting supervision to E/F/S is insufficient for accurate predictions. These considerations motivate going beyond (E/F/S) alone and integrating additional degrees of freedom and constraints to improve accuracy, stability, and transferability, which will tremendously benefit the understanding of complicated chemical systems.
To be universal, the model needs to handle both invariant and equivariant properties within a single, shared, symmetry-consistent framework, learning scalar representations for invariant quantities together with tensor representations for equivariant outputs (conserved quantities such as forces, stresses, polarization, polarizabilities; and direct properties such as direct-forces, direct-stresses, direct-dipoles, and direct-polarizabilities), so that all heads obey the expected behavior and can be co-supervised for mutual consistency.
MLIPs typically decompose the total energy into atomic site energies, which depend on the local environment of individual atoms. Early generations often relied on fixed hand-crafted representations and simple Cartesian neural networks that mapped atomic coordinates and species into invariant features. These invariant approaches achieved considerable success by ensuring translational, rotational, and permutational invariance [3, 4, 5, 6, 7]. However, they were fundamentally limited in accuracy and their ability to capture tensorial physical quantities. The advent of equivariant message passing neural networks marked a major breakthrough. By incorporating irreducible representations of the O(3) group via spherical tensors, models such as NequIP, Allegro and MACE [8, 9, 10] introduced physics-informed inductive biases that significantly improved data efficiency and generalization. This shift revolutionized the field and made spherical-tensor-based methods the mainstream choice. Nevertheless, spherical tensor formulations present intrinsic challenges. The use of Clebsch-Gordan (CG) coefficients to couple angular momenta incurs substantial computational cost (the SO(2) tensor product [11] will be discussed later). Moreover, spherical tensors are defined with respect to a fixed axis, commonly the $z$ axis, which introduces directional bias [12, 13]. In parallel, Cartesian tensor approaches have experienced a resurgence as attractive alternatives. TensorNet [14], HotPP [15], CACE [16], and CAMP [17] explored feature construction directly in Cartesian space. These models rely on reducible Cartesian tensors and aimed to improve either accuracy or computational efficiency compared to spherical-tensor-based methods. They indeed achieved some progress, particularly in low rank cases where Cartesian tensor contractions are generally faster. However, their accuracy has typically not surpassed that of models built on irreducible spherical tensor representations. TensorNet is restricted to rank-2 tensor and lacks systematic mechanisms for higher body interactions. HotPP [15] introduced deeper message passing and gating in Cartesian space but relied on reducible tensors with redundant representations. CACE encodes many body correlations explicitly but is limited in parameterization flexibility. CAMP [17] enforced strict permutation symmetry but did not guarantee tracelessness. In fact, irreducible tensors also exist in Cartesian space. Unlike spherical tensors, they do not privilege any orientation, and their tensor products are often simpler and more computationally efficient up to rank four [13]. ICTP [13], although pioneering the use of irreducible Cartesian tensors to achieve accuracies comparable to spherical tensors, requires explicit permutation of all indices in tensor products, and its implementation is limited to rank-3 tensors. Moreover, their representation cannot exploit lower-weight Cartesian tensors, thereby not fully realizing the potential of the Cartesian framework.
Here, we introduce the Tensor Atomic Cluster Expansion (TACE), a SOTA framework designed to address these challenges by unifying scalar and tensorial modeling entirely within Cartesian space. TACE systematically decomposes atomic environments into (irreducible) Cartesian tensors, ensuring exact symmetry consistency while avoiding the axis-dependence and CG overhead inherent to spherical-tensor-based methods. Building on this representation, TACE provides universal embeddings on both sides: (i) invariant embeddings for graph or node-level scalars (e.g., charges, fidelity tags, spin multiplicity) and (ii) equivariant embeddings for tensors (e.g., external fields, non-collinear magnetic moments) injected into the same feature space so that both predicted quantities and embedded quantities are handled on equal footing, enabling explicit control during the prediction process. A Latent Ewald Summation (LES) [18, 19, 20, 21] plug-in further supplies a rigorous yet efficient route to long-range interactions within the same training loop. We perform comprehensive evaluations across in-domain and out-of-domain molecular sets, condensed-phase liquids, field-perturbed solids, charged and magnetic materials, reactions systems in heterogeneous catalysis and universal MLIP benchmarks, to assess TACE’s ability to integrate diverse physical information, both simultaneously and independently, to improve prediction accuracy. Across this breadth, TACE delivers accuracy on par with or surpassing that of leading methods, while providing a systematic and extensible paradigm for modeling the complex interplay between geometry, fields, and material properties.
@@ 2 Results
Figure 1: Irreducible Cartesian tensor and the TACE architecture. a Taking the rank-2 case as an example, ICTD can extract the corresponding irreducible components with respect to a given weight. Through the basis transformation between Cartesian and spherical basis, the lower-weight parts of Cartesian tensors can be converted to $\mathrm{rank}=\mathrm{weight}$". b Illustration of the overall data flow in TACE. The embeddings of invariants and equivariants are generic. In this work we demonstrate embeddings for different computational levels, charges, external fields, and magnetic moments, but in principle other quantities are also supported. We also support the use of ZBL [22] and Latent Ewald Summation [18, 19, 20, 21] as plugins.
2.1 The TACE architecture
The feature learning of TACE begins with the construction of a graph in which each node is parameterized by learned irreducible Cartesian tensors (ICTs) ${\mathbf{h}}_{i}$ that encode its chemical environment. For each edge, a one-particle basis $\boldsymbol{\phi}_{ij}$ is computed and to obtain permutation-invariant representations of semi-local environments, a pooling operation is applied to aggregate these local features, resulting in the so-called atomic basis ${\mathbf{A}}_{i}$. Building on this, higher-body-order features are systematically constructed through tensor contraction over the ${\mathbf{A}}_{i}$, yielding a product basis ${\mathbf{B}}_{i}$ that captures ($n+1$)-body interactions. TACE follows the framework of Atomic Cluster Expansion (ACE) [23, 24, 10, 25, 26], while acknowledging that various alternative design choices remain available. In addition to the aforementioned parts, some extra features have also been added, as shown in Fig. 1b.
@@ 2.2 Cartesian tensor
A generic Cartesian tensor of rank-$\nu$, denoted as $\prescript{\nu}{}{\mathbf{T}}$, is a multilinear object with $3^{\nu}$ components defined in a three-dimensional Euclidean space. The term “generic Cartesian tensor” refers to one with no imposed symmetry or traceless constraints among its indices. These components transform according to specific representation rules under the action of the orthogonal group $\mathrm{O(3)}$, which includes both rotations and spatial inversions in three-dimensional space. In general, such tensors are reducible under $\mathrm{O(3)}$, meaning that they can be decomposed into a direct sum of irreducible components, each associated with a definite angular momentum character $\ell$. A generic Cartesian tensor $\prescript{\nu}{}{\mathbf{T}}$ can be expressed as a sum of ICTs of the same rank but different weight $\ell$ (note that this “weight” is not the “learning weight” in machine learning):
@@
where each $\prescript{{(\nu;\ell;q)}}{}{\mathbf{T}}$ is an irreducible component of rank-$\nu$, weight $l$ with $2\ell+1$ freedom, corresponding to a spherical tensor of degree $\ell$ and q denotes the label (since the multiplicity for given $\nu$ and $\ell$ may greater than 1). As a result, the generic Cartesian tensor includes all possible irreducible parts from weight $\ell=0$ up to $\ell=\nu$. This decomposition reveals that Cartesian tensors consist of a mixture of physical components: scalar parts ($\ell=0$), vector parts ($\ell=1$), and higher-rank fully symmetric-traceless tensors ($\ell\geq 2$), each transforming independently under rotation or inversion and corresponding to an invariant subspace of the tensor representation space.
@@ 2.3 Irreducible Cartesian tensor decomposition
As is well known, a generic rank-2 Cartesian tensor $\mathbf{T}$ can be decomposed into its irreducible components as follows (see also Fig. 1a):
${\mathbf{T}}_{ij}=\Bigg(\underbrace{\tfrac{1}{2}({\mathbf{T}}_{ij}+{\mathbf{T}}_{ji})-\tfrac{1}{3}\delta_{ij}{\mathbf{T}}_{kk}}_{\ell=2}\Bigg)+\Bigg(\underbrace{\tfrac{1}{2}{(\mathbf{T}}_{ij}-{\mathbf{T}}_{ji})}_{\ell=1}\Bigg)+\Bigg(\underbrace{\tfrac{1}{3}\delta_{ij}{\mathbf{T}}_{kk}}_{\ell=0}\Bigg)$ (2)
This decomposition yields the rank-2 tensor’s irreducible components corresponding to angular momenta $\ell=2$, $\ell=1$, and $\ell=0$. Generalizations to higher-rank tensors with $\nu=3,4,5$ (numerical or analytical) were proposed in 1965, 1982, and 2024, respectively [27, 28, 29]. These ICTs could be obtained via irreducible Cartesian tensor decomposition (ICTD) matrices. We denote the ICTD operater and its associated decomposition matrix as $\mathscr{T}$ [30]. Although it is possible to compute ICTs with rank $\nu>5$ using numerical methods, such approaches are not suitable for use in the forward pass of deep learning models, as they significantly degrade computational efficiency and as far as we know, there are no models utilizing numerical methods for this operation now. In fact, decomposition matrices were previously only available up to rank $\nu=5$, and their generation has traditionally been computationally expensive. However, thanks to the recent work of Shihao Shao [30], analytical ICTD matrices for $\nu>5$ can now be efficiently constructed. The overall decomposition can be written as follows, where $\operatorname{vec}(\cdot)$ denotes flattening of the Cartesian tensor and the symbol $\triangleright$ denotes an operation applied to the operand on its right-hand side.
@@
In fact, we also implemented symmetrized tensors and recursive tracelessing to obtain irreducible Cartesian tensors, which works also for arbitrary Cartesian tensors. However, this approach is ultimately purely numerical, and ICTD provides a more systematic analytical alternative.
@@ 2.3.1 Equivariant operation
The first type, tensor contracion maps two tensors $^{\nu_{1}}\mathbf{T}$ and $^{\nu_{2}}\mathbf{T}$ to a new tensor $^{\nu_{3}}\mathbf{T}$ whose rank $\nu_{3}=\nu_{1}+\nu_{2}-2k$ is determined by the the number of contracted indices $k$, where $0\leq k\leq\min(\nu_{1},\nu_{2})$. When no indices are contracted, the operation is equivalent to tensor product. Eqn. 4 gives the element-wise expression of the tensor contraction, which may be abbreviated as $\prescript{\nu_{3}}{}{\mathbf{T}}=\prescript{\nu_{1}}{}{\mathbf{T}}\boldsymbol{\otimes}_{(\nu_{1},\nu_{2},\nu_{3})}\prescript{\nu_{2}}{}{\mathbf{T}}$. In practice, certain constants such as $3^{-\tfrac{k}{2}}$ can be used to modulate the variance.
@@
We observe that for a given combination $(\nu_{1},\nu_{2},\nu_{3})$, there can be multiple distinct ways to perform contractions, each corresponding to a different contraction path. Drawing inspiration from e3nn [31], we define each such distinct contraction configuration as a path. Specifically, if $k$ indices are contracted between the first and second tensors, the number of possible paths can be computed without considering any symmetries. It equals the number of ways to select $k$ contracted indices from the first tensor and $k$ contracted indices from the second tensor, multiplied by the number of permutations of these $k$ indices. Since Cartesian tensors correspond to reducible representations and therefore contain redundant information, Cartesian-based-tensor-contraction models [15, 17] typically include only a single path for each combination and may directly embedding radial information to edge. In contrast, by employing ICTD, we have the flexibility to switch between irreducible and reducible Cartesian tensors. In the irreducible representation, the paths are often symmetric, allowing us to select only one of them, though better selection schemes may be developed in the future. We then assign a radial-dependent weight to each.
It is well known that including or not including inversion operations distinguishes the SO(3) and O(3) groups. If inversion is not considered (i.e., only SO(3) is taken into account), all spherical harmonics are assumed to have even parity within the framework of e3nn [31]. When parity is taken into account, the spherical harmonics alternate between even (e) and odd (o) parity. For models such as NequIP [8], there is an option to enable parity. Enabling parity allows the model to retain more intermediate irreducible components, but it has little effect on accuracy, which primarily depends on $\mathrm{L_{max}}$ or $\mathrm{\ell_{max}}$. Considering O(3) implies that the model internally handles polar and pseudotensors separately. In Cartesian space, the angular basis(Cartesian harmonics) in Eqn. 11 transforms in the same way as spherical harmonics. If pseudotensor prediction is required, there are two approaches:
  1. 1.
    Follow the e3nn [31] implementation and explicitly mark parity.
  2. 2.
    Convert between polar and pseudovectors by contracting tensors with the Levi-Civita tensor $\boldsymbol{\varepsilon}_{ijk}$. For example, if we want to obtain a pseudovector from two polar vectors:
$(\mathbf{\prescript{1}{}{T}_{1}}\otimes\mathbf{\prescript{1}{}{T}_{2}})_{i}=\boldsymbol{\varepsilon}_{ijk}(\mathbf{\prescript{1}{}{T}_{1}})_{j}(\mathbf{\prescript{1}{}{T}_{2}})_{k}$ (5)
The second type of operation consists of linear combinations of tensors with identical rank $\nu$, resulting in new tensors of the same rank. Specifically, given $\tilde{c}$ input tensors $\prescript{\nu}{\tilde{c}}{\mathbf{T}}$, we can construct $c$ output tensors $\prescript{\nu}{c}{\mathbf{T}}$ of the same rank through a learnable linear transformation. This operation can be interpreted as treating each tensor as a neuron, thereby enabling linear operations to be applied. In fact, similar to approaches such as NequIP and HotPP [8, 15], one could introduce nonlinearity via gating to construct multi-layer perceptrons (MLPs) for tensor with $\nu>0$.